OpenStudy (anonymous):

I WILL ABSOLUETLY GIVE MEDAL AND FAN YOU IF YOU HELP GIVE THE ANSWER TO MY THIRD LAST QUESTION PLEASE!!! A cylindrical tank has diameter 3.2 m and length 12.7 m. What is the surface area of the tank, to 1 decimal place?

OpenStudy (anonymous):

do you know what the formula for the surface area of a cylinder?

OpenStudy (anonymous):

I Just need the answer im really nervous, im gonna fail, please just answer for medal and fan

OpenStudy (jdoe0001):

hint: \(\large { \textit{total surface area of a cylinder}=2\pi r(h+r)\quad \begin{cases} h\to height\\ r\to radius \end{cases} }\)

OpenStudy (campbell_st):

you 1st need to find the radius of the cylinder then it depends on you definition of Surface Area here is Australia its all surfaces... so the formula is \[SA = 2\pi r^2 +2 \pi r h \] which is the area of the 2 circular faces... plus the rectangle with dimensions of height x circumference. hope it helps

OpenStudy (anonymous):

I don't understand! Just the answer please, I am completely running out of time!

OpenStudy (anonymous):

the formula is 2 π r ^2 + 2 π r h all you have to do is plug in the numbers that you have into the formula. so it would be 2π(1.6)^2 + 2π (1.6) (12.7) if you solve it out, it would come out to be about 143.8m^2 to the nearest tenth.